# A Decision-Language Model (DLM) for Dynamic Restless Multi-Armed Bandit Tasks in Public Health

### v3

@@ 1 Introduction
We assess our proposed method within a simulated public health environment created in partnership with ARMMAN, an India-based non-profit that spreads preventative care awareness to pregnant women and new mothers through an automated call service. To increase program listenership, ARMMAN employs health workers to provide live service calls to beneficiaries; however, due to limited support staff, ARMMAN faces a resource allocation problem to improve program listenership through optimal assignment of health workers. Prior works modeling the ARMMAN setting with RMABs have shown that RMAB policies can reduce engagement drops by $\sim 30\%$ in real-world deployment [24, 25, 26]. However, these prior works fail to address a persistent challenge within ARMMAN: tuning policies dynamically to prioritize disadvantaged groups, a common issue when targeting new regions or shifting population dynamics [27, 28]. We utilize an anonymized (Appendix B) dataset from ARMMAN to develop a simulation (Appendix C) representing changing priorities within the maternal health setting, and evaluate our method within this simulated public health environment.
@@ 2 Related Work
RMABs: The RMAB problem, introduced by Whittle [29], is classically solved through the Whittle index heuristic policy [30, 31]. Subsequent works have generalized to multi-action RMABs [32, 33]. RMABs gained prominence in public health domains and are deployed to disseminate preventative healthcare information, monitor health program adherence, and model disease spread [34, 24]. However, existing works focus on fixed reward functions and fail to consider that public health planners often have evolving priorities [3].
@@ 4.4 Reflection Stage
We propose enabling LLM self-reflection using a policy outcome comparison procedure described in Appendix D Algorithm 2. Unlike prior works [21], we do not assume access to a numerical, scalar fitness function $F$ to validate reward functions. In the absence of this scalar feedback, we propose showing state-feature distributions, described below, for LLMs to select top candidate reward functions, and use these top candidates to guide future generations. Given trained policy and critic network $\theta,\Phi$, learnt lambda value $\lambda$, and feature matrix $\mathbf{Z}\in\mathbb{R}^{N\times m}$, we first simulate over $n_{\text{s}}$ evaluation timesteps (Alg. 2 lines 5:7). For stored accumulated agent states $\mathbf{S}$, we designate preferred “positive states"; this may be customized depending on the setting, but we consider a binary per-agent state with preferred "positive" state 1. Using accumulated states, we then compute the percentage of positive states, out of all positive states observed, accrued from each key feature group $g\in\mathbf{G}$ over the evaluation timesteps, as a signal for intervention effect (Alg. 2 lines 9:11). These state-feature distributions quantify the distribution of accumulated positive states over key demographic features, ultimately describing the percentage of total positive intervention effects attributed to each discretized age range, education level, or income bracket, for example. Each distribution is reported as an output string (see Appendix L for examples).
@@ 5.1 Simulated Public Health Setting
We evaluate DLM in a simulated setting developed in collaboration with ARMMAN, an India-based non-profit that enrolls pregnant women (beneficiaries) from low-income communities [33, 24] into an automated messaging service sharing preventative healthcare information. ARMMAN employs health workers to provide live service calls to beneficiaries with a higher risk of program drop-out, however, the beneficiary population far outnumbers the available health workers. Currently, deployed RMABs in the ARMMAN setting model beneficiaries as arms, live health worker calls as intervention actions [47], and binary arm states 00 as "not engaged" and $1$ as a preferred "engaged" (positive) state when beneficiaries listen to an automated message for $>30$ seconds. The RMAB policy then identifies individuals across the entire beneficiary population to live call such that overall program engagement is maximized [48, 26]. We consider this the old, fixed “Default" policy goal. However, a key challenge for ARMMAN is training dynamic policies to support specific underrepresented groups, for example when the program is deployed in different regions [27]. We simulate these dynamic policy goals with 16 distinct prompts (Table 3) that describe demographic feature-based subpopulation priorities. We simulate this environment using a fully anonymized dataset collected by ARMMAN in January 2022, created from a service quality improvement study of 7,668 mothers; we compute transition probabilities from this dataset to simulate actions of real-world beneficiaries.
@@ 5.6 Limitations
We test our method in a purely simulated environment; any potential consideration of real-world deployment would require comprehensive field testing. Additionally, we test with prompts in English; further testing in local (Indian) languages is required. Furthermore, in any potential deployment of the proposed method, we expect that health experts should monitor state-feature distributions to intervene on proposed policies, ensuring safety and proper disambiguation of unclear input prompts.
@@ C.3 Simulated Findings and Equal Distribution of Resources
We note again that all experiments presented in this work were conducted in simulation; we attempt to simulate a plausible real-world public health setting to gain insight into strengths and weaknesses of the proposed approach. We additionally note that the intended goal of this research is to identify techniques to more easily align algorithmic resource allocation methods to specific goals of public health experts. However, the envisioned service call program is always equally accessible to all beneficiaries, and health information disseminated through the program will never be restricted for any individual enrollee. The proposed system is intended for use only in cases where separate, additional resources are available to help especially underrepresented groups. In these cases, the proposed system may help align algorithmic resource allocation of the separate additional resources more closely to desired public health expert policy goals.
@@ 35. Income bracket -1 (no income) - Binary
Appendix L Sample Reward Reflection Output
@@
L.1 Task: Older Bias
@@
L.2 Marathi Speakers + Middle Aged

### v4

@@ 1 Introduction
We assess our proposed method within a simulated public health environment created in partnership with ARMMAN, an India-based non-profit that spreads preventative care awareness to pregnant women and new mothers through an automated call service. To increase program listenership, ARMMAN employs health workers to provide live service calls to beneficiaries; however, due to limited support staff, ARMMAN faces a resource allocation problem to improve program listenership through optimal assignment of health workers. Prior works modeling the ARMMAN setting with RMABs have shown that RMAB policies can reduce engagement drops by $\sim 30\%$ in real-world deployment [24, 25, 26]. However, these prior works fail to address a persistent challenge within ARMMAN: tuning policies dynamically to prioritize disadvantaged groups, a common issue when targeting new regions or shifting population dynamics [27, 28]. We utilize an anonymized (Appendix B) dataset from ARMMAN to develop a simulation (Appendix C) representing changing priorities within the maternal health setting, and evaluate our method within this simulated public health environment. We provide a detailed discussion of our consideration of ethical guidelines during the design process of this work in Appendix A, B, C.
@@ 2 Related Work
RMABs: The RMAB problem, introduced by Whittle [29], is classically solved through the Whittle index heuristic policy [30, 31]. Subsequent works have generalized to multi-action RMABs [32, 33]. RMABs gained prominence in public health domains and are deployed to disseminate preventative healthcare information, monitor health program adherence, and model disease spread [34, 24, 35]. However, existing works focus on fixed reward functions and fail to consider that public health planners often have evolving priorities [3].
@@ 4.4 Reflection Stage
We propose enabling LLM self-reflection using a policy outcome comparison procedure described in Appendix D Algorithm 2. Unlike prior works [21], we do not assume access to a numerical, scalar fitness function $F$ to validate reward functions. In the absence of this scalar feedback, we propose showing state-feature distributions, described below, for LLMs to select top candidate reward functions, and use these top candidates to guide future generations. Given trained policy and critic network $\theta,\Phi$, learnt lambda value $\lambda$, and feature matrix $\mathbf{Z}\in\mathbb{R}^{N\times m}$, we first simulate over $n_{\text{s}}$ evaluation timesteps (Alg. 2 lines 5:7). For stored accumulated agent states $\mathbf{S}$, we designate preferred “positive states"; this may be customized depending on the setting, but we consider a binary per-agent state with preferred "positive" state 1. Using accumulated states, we then compute the percentage of positive states, out of all positive states observed, accrued from each key feature group $g\in\mathbf{G}$ over the evaluation timesteps, as a signal for intervention effect (Alg. 2 lines 9:11). These state-feature distributions quantify the distribution of accumulated positive states over key demographic features, ultimately describing the percentage of total positive intervention effects attributed to each discretized age range, education level, or income bracket, for example. Each distribution is reported as an output string (see Appendix M for examples).
@@ 5.1 Simulated Public Health Setting
We evaluate DLM in a simulated setting developed in collaboration with ARMMAN, an India-based non-profit that enrolls pregnant women (beneficiaries) from low-income communities [33, 24] into an automated messaging service sharing preventative healthcare information. ARMMAN employs health workers to provide live service calls to beneficiaries with a higher risk of program drop-out, however, the beneficiary population far outnumbers the available health workers. Currently, deployed RMABs in the ARMMAN setting model beneficiaries as arms, live health worker calls as intervention actions [48, 49, 50], and binary arm states 00 as "not engaged" and $1$ as a preferred "engaged" (positive) state when beneficiaries listen to an automated message for $>30$ seconds. The RMAB policy then identifies individuals across the entire beneficiary population to live call such that overall program engagement is maximized [51, 26]. We consider this the old, fixed “Default" policy goal. However, a key challenge for ARMMAN is training dynamic policies to support specific underrepresented groups, for example when the program is deployed in different regions [27]. We simulate these dynamic policy goals with 16 distinct prompts (Table 3) that describe demographic feature-based subpopulation priorities. We simulate this environment using a fully anonymized dataset collected by ARMMAN in January 2022, created from a service quality improvement study of 7,668 mothers; we compute transition probabilities from this dataset to simulate actions of real-world beneficiaries.
@@ 5.6 Limitations
We test our method in a purely simulated environment; any potential consideration of real-world deployment would require comprehensive field testing after approvals from relevant ethics boards. Additionally, we test with prompts in English; further testing in local (Indian) languages is required. Furthermore, whereas we present results on DLM with Chain-of-Thought (CoT) reasoning in Table 6, indicating no significant improvement due to CoT, we recognize that additional LLM reflection techniques have the potential to improve DLM. Finally, in any potential deployment of the proposed method, we expect that health experts should monitor state-feature distributions to intervene on proposed policies, ensuring safety and proper disambiguation of unclear input prompts.
5.7 Ethical Considerations in the Use of DLM
Extending our discussion of the ethical considerations we take in developing DLM (Appendix A, B, C), we further consider the broader implications of algorithmic resource allocation, particularly in public health. We note the need for mitigating data bias in the health setting, especially to minimize harmful discrimination for underrepresented groups [53], and to avoiding data bias by enabling participatory design [54], all key considerations we make in collaboration with our partner NGO for this work. We further highlight the importance of democratized decision-making criteria for algorithmic allocation techniques [55, 56], as well as the importance of complete beneficiary autonomy through guaranteed consent and the opportunity to deny allocations [57], as we do in this work. Finally, one extension of this work that future studies may consider is to directly incorporate prior work in fairness guarantees in resource allocation settings [58, 59, 60, 61].
@@ 6 Conclusion
Acknowledgments and Disclosure of Funding
This work was supported by the Harvard Data Science Initiative.
@@ C.3 Simulated Findings and Equal Distribution of Resources
We note again that all experiments presented in this work were conducted in simulation; we attempt to simulate a plausible real-world public health setting to gain insight into strengths and weaknesses of the proposed approach. We additionally note that the intended goal of this research is to identify techniques to more easily align algorithmic resource allocation methods to specific goals of public health experts. However, the service call program is always equally accessible to all beneficiaries (e.g., participants can always request service calls via a free missed call service) and health information disseminated through the program will never be restricted for any individual enrollee. That is, all participants will receive the same weekly health information by automated message regardless of whether they are scheduled to receive service calls or not. The proposed system is intended for use only in cases where separate, additional service call resources are available to help especially underrepresented groups. In these cases, the proposed system may help align algorithmic resource allocation of the separate additional resources more closely to desired public health expert policy goals.
@@ 35. Income bracket -1 (no income) - Binary
Appendix L Chain-of-Thought Experiments
We note that methods that improve LLM reasoning, such as chain-of-thought (CoT), may be incorporated to improve the base LLM reasoning in the DLM pipeline. Below, we include results using chain-of-thought reasoning for our model outputs; while we did not find that CoT significantly improved reward, we nevertheless emphasize the potential for additional LLM techniques to improve DLM outputs.
Table 6: Full prompts and numerical results with chain-of-thought (CoT) reasoning for DLM.
idx Label Full Prompt Base (topline) No Action Default DLM with CoT (No Reflection) DLM with CoT (Reflection)
0 Older Bias While still prioritizing all, slightly focus on the oldest by age distribution. 1.00 -0.02 ± 0.072 0.41 ± 0.130 0.89 ± 0.034 0.95 ± 0.023
1 Low Income While still prioritizing all, slightly focus on the bottom 12.5% of the income_bracket distribution. 1.00 -0.02 ± 0.073 0.51 ± 0.122 0.71 ± 0.062 0.62 ± 0.084
2 Hindi Speakers While still prioritizing all, slightly focus on those who speak Hindi. 1.00 -0.12 ± 0.067 0.71 ± 0.093 0.79 ± 0.055 0.94 ± 0.032
3 Low Education Bias While still prioritizing all, slightly weight those who have had low education. 1.00 -0.00 ± 0.073 0.46 ± 0.124 0.79 ± 0.059 0.72 ± 0.065
4 Age Distribution Tail Emphasis While still prioritizing all, slightly focus on both the youngest and oldest by age. 1.00 -0.02 ± 0.069 0.52 ± 0.111 0.64 ± 0.062 0.78 ± 0.044
5 Middle Income While still prioritizing all, slightly prefer the income bracket bounds for the middle 40% of the population. 1.00 -0.08 ± 0.067 0.65 ± 0.097 0.69 ± 0.072 0.66 ± 0.083
6 Non-Phone Owner While still prioritizing all, slightly favor those women who do not own their own phone. 1.00 -0.07 ± 0.067 0.50 ± 0.114 0.92 ± 0.029 0.85 ± 0.038
7 Low Income + Younger Bias While still prioritizing all, slightly prioritize impoverished younger mothers by combining the distributions of ’age’ and ’education’. 1.00 0.03 ± 0.075 0.40 ± 0.134 0.91 ± 0.021 0.96 ± 0.019
8 Marathi Speakers + Middle Aged While still prioritizing all, slightly focus on those Marathi-speakers with middle-aged mothers. 1.00 -0.00 ± 0.074 0.51 ± 0.115 0.88 ± 0.033 0.91 ± 0.033
9 Early and Late Workers While still prioritizing all, slightly emphasize beneficiaries who likely work early in the morning and late at night. 1.00 -0.11 ± 0.065 0.69 ± 0.093 0.93 ± 0.035 0.94 ± 0.024
10 Critical Low Income While still prioritizing all, slightly weight the lowest income_bracket groups, the absolute lowest earners in the population. 1.00 -0.08 ± 0.068 0.60 ± 0.104 0.85 ± 0.041 0.83 ± 0.041
11 Early Morning Call + NGO Registered While still prioritizing all, slightly advantage those who prefer being called before 10:30am ’slot’ and are registered at an NGO. 1.00 -0.12 ± 0.065 0.70 ± 0.085 0.94 ± 0.027 0.94 ± 0.027
12 Morning Call + NGO Registered While still prioritizing all, slightly advantage those who prefer being called between 10:30am-12:30pm and are registered at an NGO. 1.00 -0.07 ± 0.066 0.52 ± 0.109 0.93 ± 0.026 0.92 ± 0.031
13 Afternoon Call + NGO Registered While still prioritizing all, slightly advantage those who prefer being called between 12:30pm-3:30pm and are registered at an NGO. 1.00 -0.11 ± 0.062 0.72 ± 0.085 0.76 ± 0.052 0.82 ± 0.041
14 Evening Call + NGO Registered While still prioritizing all, slightly advantage those who prefer being called after 7PM ’slot’ registered at an NGO. 1.00 -0.06 ± 0.070 0.55 ± 0.111 0.83 ± 0.046 0.78 ± 0.045
15 Technically Challenged While still prioritizing all, infer technical challenges in reaching the phone that could indicate ’at-risk’ beneficiaries and give slight preference. 1.00 -0.14 ± 0.060 0.64 ± 0.097 0.80 ± 0.045 0.90 ± 0.037
Appendix M Sample Reward Reflection Output
@@
M.1 Task: Older Bias
@@
M.2 Marathi Speakers + Middle Aged
@@
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction clearly state and claims, and the claimed made in the abstract and introduction match the theoretical and experimental results.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The scope of the claims are clearly stated (see abstract, introduction, and experiments). Our approach is designed for restless multi-arm bandits tasks, and the experiments are focused on public health domains.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: We provide all details on theoretical results, including necessary assumptions and detailed derivations.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We disclose all information needed to reproduce the experimental results, including training details (Section 5.3), a full task list and ground truth base reward function (see Appendix E), full output of sample reward reflection procedure (see Appendix F).
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [No]
  24. Justification: We use a real world dataset collected by a non-profit organization in public health domains. We are not allowed to release the data. However, we will release the code, and we describe dataset details (e.g. dataset size, features available) and dataset consent for collection and analysis in Appendix B and  C.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: In Section 5 on experiments, we provide detailed of the experimental setting, including details of the simulated public health setting, tasks, baselines, metrics, training details, and evaluation details.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We report error bars in the main experimental results (see Figure 2). We report how these errors bars are calculated (number of trails, number of samples per steps, etc) in Figure 2 caption. We also report results on statistical significance, specifically one-tailed t-test results, in Table 5 in Appendix.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We report information on computing resources in Appendix Section E
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Our experiments confirm with NeurIPS code of ethics. For experiments on real-world data, we describe dataset consent for collection and analysis in Appendix C.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We discuss societal impacts of this work in the social impact statement (see Section A in Appendix).
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [Yes]
  54. Justification: We discuss safeguards in place for responsible and safe use of data and models in Section 5. The public health information dataset is completely anonymized by the non-profit organization we collaborate with (see Section C for more details).
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We cite the original papers that produce the code and models used. Specifically, in Section 5.3, we clearly state that we use the Gemini Pro LLM model [1] from Google to generate reward functions, and train our downstream policy with RL using PPO [52].
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: This paper does not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This paper does not involve crowdsourcing nor research with human subjects. We conduct a secondary analysis of the ARMMAN dataset (see Appendix C for details).
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [Yes]
  74. Justification: The data exchange from ARMMAN to the researchers was regulated through clearly defined exchange protocols including anonymization, read-only researcher access, restricted use of data for research purposes only, and approval by the ARMMAN ethics review committee (see Appendix C for details).
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.